Solve
2x/3-x=x/6-3/2

Answer by addingup(3677)

Here's two options, depending on what you meant by 2x/3-x:
(2x/3) - x = (x/6) - (3/2); x = 3
(2x)/(3x) = (x/6) - (3/2); x = 13